OpenStudy (hhopke):

Okay... I have two questions. What is the factored form of -x^2-6x=0? And what is the factored form of -4x^2=8x?

OpenStudy (gabebae):

Hello I'd love to help you there

OpenStudy (anonymous):

-x(x +6) = 0 -4x(x +2) = 0

OpenStudy (hhopke):

Okay. Thanks, @someone7070! And thanks for the offer, @GabeBae

OpenStudy (gabebae):

I was a bit afk I was waiting for you to respond